OpenStudy (anonymous):

-2 1/2 divided by 6?

OpenStudy (123ab456c):

why dont you just use a calculator

OpenStudy (anonymous):

I'd like to know how to do it since I keep getting the wrong answer

OpenStudy (anonymous):

\[-2\frac{ 1 }{ 2 }=\frac{ -3 }{ 2 }\div6=\frac{ -3 }{ 2 }\times \frac{ 1 }{ 6 }\] hope that helps

OpenStudy (123ab456c):

the answer is 5/12

OpenStudy (anonymous):

- 5/12?

OpenStudy (123ab456c):

no just 5/12

OpenStudy (anonymous):

No it's negative 5/12

OpenStudy (anonymous):

It's a negative divided by a positive that means it is negative

OpenStudy (123ab456c):

just becuase it has a negative dosent mean it will come out as one

OpenStudy (anonymous):

Ummm yes it does especially if the other term isn't negative

OpenStudy (anonymous):

But isn't the rule : negative divided by positive = negative?

OpenStudy (anonymous):

Type it into a calculator Yes it is @LoveIt

OpenStudy (anonymous):

ok

OpenStudy (123ab456c):

oh wait it is is so its -5/12

OpenStudy (123ab456c):

your welcome for telling you the answer!

OpenStudy (anonymous):

thanks.

OpenStudy (123ab456c):

can you be my fan and give me a medal plese:)